) For a consumer, the marginal utility of good A is 25 and its price is $5. The marginal utility of good B is 60 and its price is $12. The consumer has allocated his entire budget. Is this consumer maximizing his total utility? Explain your answer.

b) "If Bill's income increases from $80,000 per year to $91,000 per year and he consumes pickup trucks and lamb chops with his money, Bill's budget line shifts outward and the increase in income means he can consume more trucks and more lamb." Is this statement true or false? Briefly explain your answer.

c) What are diseconomies of scale and why might they occur?

d) Martha's Cleaning Services is a perfectly competitive firm that currently cleans 30 offices a week and charges $20 per office, which is the going market price. Martha's marginal cost is $15. What should Martha do to increase her economic profit? Clean more offices? Raise her price? Explain your answer.
1
Expert's answer
2020-06-18T13:49:22-0400

For a consumer, the marginal utility of good A is 25 and its price is $5. The marginal utility of good B is 60 and its price is $12. The consumer has allocated his entire budget. Is this consumer maximizing his total utility? Explain your answer.


The consumer will maximize her utility at the point where:



"\dfrac{MU_A}{MU_B} = \dfrac{P_A}{P_B}"

In our question, we have


"MU_A = 25, \; P_A = \$5\\[0.3cm]
MU_B = 60,\; P_B =\$12"


Therefore:



"\dfrac{25}{60} = \dfrac{5}{12}\\[0.3cm]
\dfrac{5}{12} = \dfrac{5}{12}"

Since the condition for utility maximization is satisfied, the consumer is maximizing her utility.



b) "If Bill's income increases from $80,000 per year to $91,000 per year and he consumes pickup trucks and lamb chops with his money, Bill's budget line shifts outward and the increase in income means he can consume more trucks and more lamb." Is this statement true or false? Briefly explain your answer.


The statement is true. An increase in the consumer's income with prices held constant shifts the budget line outwards, meaning that more units of each good is consumed.



c) What are diseconomies of scale and why might they occur?


Diseconomies of scale are the cost disadvantages that a firm experiences due to its increased scale of production. It happens when the firm's average total cost increases as the output increases.


d) Martha's Cleaning Services is a perfectly competitive firm that currently cleans 30 offices a week and charges $20 per office, which is the going market price. Martha's marginal cost is $15. What should Martha do to increase her economic profit? Clean more offices? Raise her price? Explain your answer.


Since she is operating in a perfectly competitive market, she should opt offering her cleaning services at the point where "P = MC" . Given the information, "P = \$20>MC = \$15" . This means that Martha is producing below the optimal point and she can do better by increasing the cleaning services so that the marginal cost increases until it is equated to the marginal cost.